OpenStudy (verifiednykia):

if tan θ = 12/5 and cos θ<0, what is sin θ?

OpenStudy (verifiednykia):

these are the choices https://gyazo.com/28d732e17e4f4c5ffd4ee0233aa34194

OpenStudy (michele_laino):

hint: using data of the exercise, we have the subsequent drawing:  namely the subsequent condition holds: \(\pi < \theta <3 \pi/2\)

OpenStudy (michele_laino):

now we have to apply this condition: \[{\left( {\cos \theta } \right)^2} + {\left( {\sin \theta } \right)^2} = 1\] where: \[\sin \theta = \frac{5}{{12}}\cos \theta \] please replace \(\sin \theta\) from second formula into the first one, and solve for \(\cos \theta\)

OpenStudy (michele_laino):

oops.. the second formula, is: \[\sin \theta = \frac{{12}}{5}\cos \theta \]

OpenStudy (michele_laino):

and here is the right drawing: